OpenStudy (anonymous):

evaluate. (x)2 - 3x, for x = -5

OpenStudy (anonymous):

-10 - -15 = 5 -10 + 15 = 5 subtracting a negative number means adding it

OpenStudy (anonymous):

If x = -5, then substitute -5 for all x's appearing in the statement. So (x)2 - 3x translates to (-5)*2 - 3*(-5) = (-10) - (-15) = (-10) + 15 = 5.

OpenStudy (anonymous):

thank i understand now